The Coronavirus disease has led to many deaths globally. In Gauteng, the number of deaths per week increases by 9.2% every week. Suppose this trend does not change and 82 deaths have been recorded this week. Determine the number of deaths that: a) were recorded 4 weeks ago b) will be recorded in 3 weeks’ time c) will be recorded in 6 weeks’ time [1] a) 63 b) 98 c) 127 [2] a) 58 b) 107 c) 139 [3] a) 63 b) 106 c) 138 [4] a) 58 b) 108 c) 127


1
Expert's answer
2022-02-03T11:31:47-0500

The number of deaths that:

a) were recorded 4 weeks ago is:

"82\/1.092^4 = 57.66;"

b) will be recorded in 3 weeks’ time is:

"82\u00d71.092^3 = 106.77;"

c) will be recorded in 6 weeks’ time is:

"82\u00d71.092^6 = 139.04."

So, the correct answer is [2] a) 58 b) 107 c) 139.